# Stereotype or Personalization? User Identity Biases Chatbot Recommendations

### v1

@@ Abstract
We demonstrate that when people use large language models (LLMs) to generate recommendations, the LLMs produce responses that reflect both what the user wants and who the user is. While personalized recommendations are often desired by users, it can be difficult in practice to distinguish cases of bias from cases of personalization: we find that models generate racially stereotypical recommendations regardless of whether the user revealed their identity intentionally through explicit indications or unintentionally through implicit cues. We argue that chatbots ought to transparently indicate when recommendations are influenced by a user’s revealed identity characteristics, but observe that they currently fail to do so. Our experiments show that even though a user’s revealed identity significantly influences model recommendations ($p<0.001$), model responses obfuscate this fact in response to user queries. This bias and lack of transparency occurs consistently across multiple popular consumer LLMs (gpt-4o-mini, gpt-4-turbo, llama-3-70B, and claude-3.5) and for four American racial groups. 111 All code and data to reproduce analyses will be released with the final version of the paper.
@@ 1 Introduction
Figure 1: A user asks for options with different levels of identity disclosure. Indicating that they are Black results in biased responses, whether intended [ ] or not [ ]. When asked, the response obfuscates the impact of race.
@@
We posit that this mechanism embeds pragmatics into conversations Grundy (2019) between people and LLMs. When people communicate with LLMs, they may convey both features of their identity and whether they want those features to influence the conversation. For example, in Figure 1, we see three ways in which a high school student might ask for advice on what universities to consider. The baseline form of the question does not reveal identity, so the system’s default assumptions play a larger part in its recommendations. In this case of Standard American English (SAE), models recommend prestigious institutions in the United States – which, historically, have admitted mostly white students. On the other hand, when the user includes explicit indications that they are Black, the system diverts to recommending historically Black colleges and universities (HBCUs).333 These are historically Black US colleges and universities, founded prior to 1964, “whose principal mission was, and is, the education of black Americans” U.S. Code (1965)
However, in cases where identity is revealed implicitly and unintentionally, recommendations may be undesirably biased toward stereotypes. In the final example, the user does not explicitly mention their race, but they mention volunteering with the National Urban League, an organization that primarily serves black communities in major cities. While the user gives no indication that they want the inferred race to inform the recommendations, the system generates recommendations informed by assumptions on the user’s identity, and risks stereotyping the user based on unintentionally shared characteristics.
Furthermore, the system obfuscates the influence of identity in generating the recommendations. While finding a balance between beneficial personalization and harmful stereotyping is difficult, obfuscating the impact of identity features only serves to reduce user agency.
@@ 3. RQ3:
We generate synthetic recommendation requests covering a variety of user identities, signals, and constraints. Aggregating LLM responses to these requests, we measure the real-world demographic alignment of the recommendations to the user’s identity using sources such as he U.S. Department of Education’s College Scorecard U.S. Department of Education (2024) and the American Community Survey U.S. Census Bureau (2016). Ultimately we find an implicit identity effect across all models tested: models generate biased recommendations regardless of the user’s expressed desire for personalization, and subsequent responses will obfuscate this bias from the user, reducing user agency and reinforcing existing stereotypes.
@@ 2 Methodology
To study the degree of bias in LLM-generated recommendations, we send LLMs prompts that request recommendations, comparing results across three levels of identity disclosure: implicitly, explicitly, or not at all. We then measure the real-world demographic distribution associated with the recommended entities, to determine whether the model produces recommendations that demographically correlate with the user’s revealed identity. Figure 2 illustrates the overall process.
@@ 2.1 Identity Selection
While bias can be examined across many axes of identity (socioeconomic class, gender, sexuality, etc.), in our analysis we specifically focus on race within the context of the United States for three reasons. First, demographic data on race is collected by the U.S. government and is widely available for a variety of things that may be recommended, such as universities U.S. Department of Education (2024) and neighborhoods U.S. Census Bureau (2016). Second, there is well-documented linguistic variation associated with racial groups in the form of different dialects of American English. These variations allow us to provide implicit signals linked to a particular racial identity. Last, we aim to study whether LLMs regurgitate decades of deeply ingrained and noted racial discrimination in American society.
@@ 2.2 Constructing Prompts
We develop an automatic procedure for composing realistic recommendation request prompts. Each prompt contains user requests with optional constraints and either explicit or implicit identity disclosures. For implicit identity disclosures, we consider implicitly revealed identity through either reference to demographically associated entities or the use of demographically linked dialects.
We set up each prompt as a concatenation of sentences, each of which expresses one of the preceding prompt components. Figure 2 “Prompt Generation” demonstrates a simplified view of how many variants can be created for each request, covering a range of user desiderata, biographical details, and revealed identities.
Explicit disclosures
Implicit disclosures
Implicit disclosures are the most complex and can come in the form of either references to stereotypical, demographically associated entities or the inclusion of dialectal features. Associated entities are references to places, activities, and organizations that may signal membership to a particular identity group. Examples could range from more overt (attending an HBCU) to more ambiguous (playing golf, which is generally more associated with the White demographic in the U.S.) Closson and Henry (2008); Davidson (1982).
Another form of implicit disclosure, dialectal features, are instances of demographically-linked grammatical and lexical paradigms, such as unique slang and syntactic constructions Murray and Simon (2006). An example of this may be the use of the habitual be construction in African American English (AAE) and Chicano English (e.g. I be in school instead of I am in school) Schneider (2008). We select these features from linguistic resources on American dialects Schneider (2008); Hanna (1997). We use Standard American English Trudgill (2002) as taught in schools as the baseline due to most dialectal speakers having the ability to code-switch into it in more formal contexts DeBose (1992). Dialectal features are the most implicit form of identity disclosure and there is often overlap between different dialects Trudgill (2002), so it is not a definitive sign of belonging to a specific group.
Requests and constraints
@@ 2.3 Measuring Demographic Alignment
Following each request, we calculate the real-world demographic alignment of each recommendation to the revealed identity of the request. We approach this in a few ways. Primarily, we examine the percentage share of people of the user’s race in each recommended university or neighborhood, pulled from U.S. government statistics, then average across the recommendations provided. For a demographic group $g$ with recommendations $R$:
@@
Specifically for universities, we determine whether an LLM recommends a wider or narrower set of options than are reflected in real-world statistics. We limit this analysis to universities because most students only attend a single undergraduate institution, making it easy to calculate outcomes. Collecting all recommended universities $R^{\prime}$ occurring in recommendations produced for users of demographic group $g$, we compute the proportion of U.S. undergraduate students belonging to group $g$ that are covered by the recommendations:
$\text{Coverage}(g)=\frac{1}{N}\sum_{r\in R^{\prime}}\text{\# of }g\text{ people in }r$
The total number of students belonging to group $g$, denoted $N$, is based on the total undergraduate student population of the United States, reported by the U.S. Department of Education U.S. Department of Education (2024).
@@ 2.4 Experimental Design
To align the racial groups with documented dialectal and data boundaries, we focus on four categories: White, Black, Hispanic, and Asian.444 Throughout this paper, we capitalize in accordance with the APA Style Guide on racial identity. Dialect features were chosen in accordance with the prompt’s syntactic structure manually from linguistic resources Trudgill (2002); Hanna (1997). We examine the results on a range of popular consumer chatbots555 Models were prompted in August 2024.: gpt-4o-mini, gpt-4-turbo, llama-3-70b-instruct-lite, and claude-3.5-sonnet-20240620. We evaluate on two tasks: university and neighborhood selection. For universities, we focus our scope on American universities but otherwise leave the location unspecified. For neighborhoods, we specify one of four cities: New York City, Los Angeles, Boston, and Chicago.
To identify a list of entities that an LLM associates with each demographic group, each model we evaluated was prompted with “What are 10 <entity type> associated with <race> people?” Entities that appeared in the top-10 list of only one race were incorporated into the list of associated entities for that race. We treat any entities that appeared in multiple lists (e.g. “the YMCA”) or as non-specific forms (e.g. “a club”) as neutral entities. Associated entities can be more or less revealing of identity, but are not definitive signals of identity even if certain racial groups are stereotypically associated with them (i.e. non-Asian students may join a Japanese film club).
Due to the combinatorial complexity of the number of possible queries ($n>10!$), we instead sample 10,000 queries from the set of all possible combinations of identity signals and constraints for each race and model, for a total of 160,000 synthetic user requests. The template for the prompts we used can be found in Appendix A. To identify the real-world demographics for each recommendation, we use official US government sources. For university and neighborhood racial demographics, this was relatively straightforward, being pulled from the Department of Education’s College Scorecard U.S. Department of Education (2024), the 2010 U.S. Census U.S. Census Bureau (2010) and the American Community Survey U.S. Census Bureau (2016).
@@ 3 RQ1: Inferred Identity Biases Recommendations
Figure 3: Mean share of each race for the recommendations provided by all evaluated models across the inclusion of dialectal features [D], associated entities [A], and explicit indications [E]. Left boxplots [ ] represents users of the same race as the figure’s calculation and right boxplots [ ] represents users of all other races. Shaded range [ ] is the range (25th - 75th percentile) of race share for baseline prompts across models.
In our first experiment, we examine our first research question: whether LLMs provide recommendations that are biased towards the perceived racial identity of the user, regardless of whether identity is explicitly revealed.
@@ 3.1 Demographic Alignment by Signal
A model that personalizes only when the user explicitly asks for their racial identity to factor into decisions – and otherwise avoids incorporating stereotypes – should behave differently depending on the type of identity disclosure. When identity is revealed implicitly, with unclear pragmatics, the demographic alignment (in this case, the percentage of people associated with that entity of the user’s race) we might expect recommendations that are no different from those where identity is revealed at all. On the other hand, when identity is indicated explicitly, we expect models to favor recommendations that have greater demographic alignment.
Figure 3 shows the degree of demographic alignment for each user identity averaged across models. We see that models respond strongly to explicit signals of user identity by providing recommendations that have comparatively high shares of the user’s race. For example, in the plot of Black users’ results for neighborhood selection, we see that when a user requests recommendations for neighborhoods and includes “keep in mind, I am Black," they receive recommended neighborhoods with around $18.5\%$ more Black residents on average than a user who does not reveal their identity. This personalization based on explicit identity disclosure is seen for users of all races except White users, who find their results mostly still fall squarely within the range of those given to baseline users.
We do see varying numerical differences in alignment across races, with Black users seeing a much starker jump than Asian users, but this is likely due to the broader demographic makeup of the United States. Unlike for Black students, there are very few (if any) institutions that are majority Hispanic or Asian, so the upper bound on the average share of their races is lower.
Models also produce demographically aligned results when users reveal their identity implicitly, either by referencing stereotypically associated entities (i.e. referencing their undergraduate degree at an HBCU), or when a user writes their request in a racially-associated dialect. This occurs to a lesser extent than it does with explicitly disclosed identity, but the difference is still statistically significant when the user references associated entities. In the context of important life decisions like where to go to school or where to live, basing decisions on unintentionally disclosed racial identities can perpetuate stereotyping without the knowledge of the user. Additionally, there is the risk of error; not everyone who references an entity belongs to the stereotypically associated race (e.g., anyone can play basketball) and not everyone who speaks a specific racial dialect identifies with that race, especially with the internet offering more cross-cultural interactions Reyes (2005); Roth-Gordon et al. (2020).
@@ 3.2 Recommendation Diversity
Figure 4: Percentage of total US students of each race represented by the university recommendations for members of that race, across all model samples. Total number of recommendations is the same across races.
Besides the demographic alignment of recommendations, there is a disparity in the diversity of options provided to users of different races. In theory, across a large number of prompts where students state varying preferences and constraints, the racial distribution of the aggregate set of recommended universities should approximate the racial distribution of undergraduate students in the United States as a whole. Put simply, a model should not artificially restrict users to only the most stereotypical recommendations for their racial group, even if the model is attempting to personalize results. If this were to be the case, we would expect see the same percentage of real-life student outcomes (the institutions where students of that race ended up) covered for all four racial groups.
In Figure 4, we see that this is not the case. Over $50\%$ of White students’ real-life outcomes are covered across the recommended universities, but less than $35\%$ of students’ outcomes are covered for each of the other races. This indicates that once a user expresses their race, either explicitly or implicitly, users from minority backgrounds receive more highly concentrated recommendations than White users do, resulting in a relatively narrower set of overall options for minority users.
@@ 3.3 Stereotypical Descriptions
White Black Hispanic Asian
journalism activism serving music
historic woman language ice
larger continued affiliation performance
players color translation international
suburban empowerment related relations
@@
White Black Hispanic Asian
extremely historically majority influenced
studios biking roots ethnicities
trail safety airport enclave
rates safer inspired markets
boutiques development beautiful significant
@@
Table 1: Words used to explain recommendations across models for dialect-only samples. Words common to specific demographics are identified using pointwise mutual information (PMI) and Kneser–Ney smoothing. We exclude words ocurring fewer than 10 times.
Even beyond the recommendations themselves, we find language models bias the values they latently attribute to each user, represented by the differences in how the model justifies each recommendation it makes. To examine this, we find the words that have the highest association with a particular demographic group using pointwise mutual information (PMI) for each word $w$ present in the explanation for group $g$ (after removing stopwords, words that appear in the query, and names of recommended entities), calculating only for words that appear more than 1000 times across the 800,000 generated explanations:
$\text{PMI}(w,g)=\log_{w}\frac{P(w|g)}{P(w)}$
Because we are ranking words for each group, and because there are a fixed number of prompts and responses, we can simplify the formula:
$\text{Score}(w,g)=\frac{\text{\# of }w\text{ in }g\text{ responses }}{\text{\#of }w\text{ in all responses }}$
We omitted samples with references to associated entities that included the names of racial groups to avoid extra bias from those racial terms (i.e. cuisines such as “Italian”). Additionally, we applied Kneser–Ney smoothing Ney et al. (1994) to reduce the impact of rare words that occur only once in the dataset.
In Table 1, we see that there are different terms used to describe recommendations for each race. For example, White students seeking university recommendations are recommended options explained by their historic past, suburban setting, or size, while Black, Hispanic, and Asian students see models heavily focusing on their perceived race and culture: Black students see more emphasis on activism and empowerment, Hispanic students see an emphasis on language and institutions being minority-serving, and Asian students see a focus on music, performance, and an international aspect. These findings occur across prompts with similar academic scores and preferences (as we indicated these to the model through sampling constraints and biographical information). By generating explanations that assign different values and preferences to different racial groups, our study reveals that models not only produce biased recommendations, but that the generated rationales reflect (and possibly reinforce) the stereotypes that already exist in American culture.
Likewise, different groups are assumed by the model to have different unexpressed preferences for locations when selecting a neighborhood. White users see an emphasis on studios and boutiques. In contrast, Black users see an outsized focus on safety despite an equal number of users across races indicating a preference for safer neighborhoods in our prompts. Additionally, Asian users are recommended neighborhoods described by the response as ethnic enclaves, which hearkens back to the history of redlining in American cities where the "co-occurrence of other social conditions linked to poverty" disproportionately impacted racial minorities living in these enclaves a "result of decades of racism and disinvestment in communities of color" Lynch et al. (2021). Again, by focusing on stereotypical attributes for non-White user prompts while providing White users with recommendations for neighborhoods with features implying affluence (i.e. boutiques), these models may perpetuate the enduring discriminatory effects of historical segregation, especially when deployed at scale.

### v2

@@ Abstract
While personalized recommendations are often desired by users, it can be difficult in practice to distinguish cases of bias from cases of personalization: we find that models generate racially stereotypical recommendations regardless of whether the user revealed their identity intentionally through explicit indications or unintentionally through implicit cues. We demonstrate that when people use large language models (LLMs) to generate recommendations, the LLMs produce responses that reflect both what the user wants and who the user is. We argue that chatbots ought to transparently indicate when recommendations are influenced by a user’s revealed identity characteristics, but observe that they currently fail to do so. Our experiments show that even though a user’s revealed identity significantly influences model recommendations ($p<0.001$), model responses obfuscate this fact in response to user queries. This bias and lack of transparency occurs consistently across multiple popular consumer LLMs and for four American racial groups.  [^1] https://github.com/AnjaliRuban/llm-stereotype-or-personalization
[…]
Stereotype or Personalization?
User Identity Biases Chatbot Recommendations
Anjali Kantharubanthanks:  Equal Contribution  Jeremiah Milbauer Maarten Sap  Emma Strubell  Graham Neubig Carnegie Mellon University anjaliruban@cmu.edujmilbaue@andrew.cmu.edu msap2@andrew.cmu.eduestrubel@cs.cmu.edugneubig@cs.cmu.edu
@@ 1 Introduction
Figure 1: Asking for options with different levels of identity disclosure. Indicating that they are Black results in biased responses, whether intended [ ] or not [ ]. When asked, the response obfuscates the impact of race.
@@
We posit that this mechanism embeds pragmatics into conversations Grundy (2019) between people and LLMs. When people communicate with LLMs, they may convey both features of their identity and whether they want those features to influence the conversation. For example, in Figure 1, we see three ways in which a high school student might ask for advice on what universities to consider. The baseline form of the question does not reveal identity, so the system’s default assumptions play a larger part in its recommendations, recommending prestigious institutions that, historically, mostly admitted white students. On the other hand, when the user includes explicit indications that they are Black, the system recommends historically Black colleges and universities (HBCUs).333 These are historically Black US colleges and universities, founded prior to 1964, “whose principal mission was, and is, the education of black Americans” U.S. Code (1965) In the final example, the user does not explicitly mention their race, but they mention volunteering with an organization that primarily serves black communities in major cities. The user gives no indication that they want the inferred race to inform the recommendations, but the system generates recommendations informed by its assumptions, and risks stereotyping the user.
Furthermore, the system obfuscates the influence of race in generating the recommendations. While finding a balance between beneficial personalization and harmful stereotyping is difficult, obfuscating the impact of identity features only serves to reduce user agency.
@@ 3. RQ3:
We generate synthetic recommendation requests covering a variety of user identities, identity signals, and constraints. Aggregating LLM responses to these requests, we measure the real-world demographic alignment of the recommendations to the user’s identity using official US government sources. Ultimately we find an implicit identity effect across all models tested: models generate biased recommendations regardless of the user’s expressed desire for personalization, and subsequent responses will obfuscate this bias from the user, reducing user agency and reinforcing existing stereotypes.
@@ 2 Methodology
To study the degree of bias in LLM-generated recommendations, we compare their responses across three levels of identity disclosure: implicit, explicit, or none. We then measure the real-world demographic distribution associated with the recommended entities to determine whether the model produces recommendations that are demographically biased toward the user’s revealed identity. Figure 2 illustrates the process.
@@ 2.1 Identity Selection
Although bias can be examined on many axes of identity (socioeconomic class, gender, sexuality, etc.), in our analysis we specifically focus on race within the United States for three reasons. First, the U.S. government collects demographic data on race for a variety of entities of interest, such as universities U.S. Department of Education (2024) and neighborhoods U.S. Census Bureau (2016). Second, there are distinctive linguistic variations associated with English-speaking racial groups in the U.S. that have been well documented Schneider (2008). Last, the systematic and ingrained racial discrimination in American society has been heavily studied, and identifying LLM regurgitation of these themes is of great importance, and has been the focus of significant previous work in other contexts Caliskan et al. (2017); Garg et al. (2018); Hofmann et al. (2024a).
@@ 2.2 Constructing Prompts
We develop an automatic procedure for composing realistic recommendation request prompts. Past work has used synthetic queries for the purposes of isolating sources of bias, since natural queries often introduce multiple degrees of variation from one another Wan et al. (2023a); Röttger et al. (2024); Castricato et al. (2025). Our prompts are more complex than just simple templates and they are modeled on the way people actually interact with models Wang et al. (2024) – rather than asking models blunt survey questions designed for humans Tjuatja et al. (2024). These synthetically generated prompts allow us to have much greater coverage of diversity identities and to systematically vary features to derive much more rigorous and sound results than if we sourced naturally written human prompts. Each prompt contains user requests with optional constraints and either explicit or implicit identity disclosures. For implicit identity disclosures, we consider implicitly revealed identity through either reference to demographically associated entities or the use of demographically linked dialects.
These implicit identity disclosures, both associated entities and dialectal features, were chosen by asking each model to generate a list of racially associated entities within a category. The lists were then manually compressed to remove items that occurred across lists for multiple races, items that were not attested in linguistic resources (for dialectal features) and items that explicitly mention the race of the user. More information about prompt generation can be found in Appendix B.
We use Standard American English Trudgill (2002) as taught in schools as the dialectal baseline due to most dialectal speakers having the ability to code-switch into it in more formal contexts DeBose (1992). For associated entities, we use generic references ("I play sports" as opposed to "I play football") or entities that appeared across the generated lists for multiple races.
Last, we provided non-racial constraints and requests as a method of diversifying the model generations. Examples include varying the standardized testing scores, budgets, and preferences of the user. This ensures that notions such as perceived socioeconomic class and ability play a lesser role.
@@ 2.3 Measuring Demographic Alignment
Following each request, we calculate the real-world demographic alignment of each recommendation to the revealed identity of the request. Information on matching recommendations with entities reported on can be found in Appendix C.1. We approach alignment in a few ways. Primarily, we examine the percentage share of people of the user’s race in each recommended university or neighborhood, pulled from U.S. government statistics, then average across the recommendations provided. For a demographic group $g$ with recommendations $R$:
@@
Additionally, we investigate the diversity and representativeness of recommendations in the best case where personalization is appropriate – when users pragmatically indicate to do so through explicit disclosure. We compare recommendation distributions with real-world data on neighborhood population (where data is more readily available).
@@ 2.4 Experimental Design
To align the racial groups with documented dialectal and data boundaries, we focus on four categories: White, Black, Hispanic, and Asian.444 Throughout this paper, we capitalize in accordance with the APA Style Guide on racial identity. Dialect features were chosen in accordance with the prompt’s syntactic structure manually from linguistic resources Trudgill (2002); Hanna (1997). We examine the results on a range of popular consumer chatbots: GPT 4o Mini, GPT 4 Turbo, Claude 3.5 Haiku, Claude 3.5 Sonnet, Llama 3.1 70B, Llama 3.1 405B, and Gemini 1.5 Pro. Exact specifications on model configuration can be found in Appendix A.
We evaluate on two tasks: university and neighborhood selection. For universities, we focus our scope on American universities but otherwise leave the location unspecified. For neighborhoods, we specify one of three cities: New York City, Los Angeles, and Chicago. Prompting details, and the exact criteria and contents for the group-specific and baseline entity sets, are in Appendix B, as well as details on how we validate our sampling setup in Appendix E.
Due to the number of possible queries ($n>10!$), we sample $8,600$ queries for each model for both university and neighborhood recommendations, for a total of 120,000 synthetic user requests. The template for the prompts we used can be found in Appendix B. We constrain the outputs to be in a JSON format for analysis, but we find that the trends we see hold in the unconstrained setting in Appendix E. We align recommendations with demographic data from official US government sources U.S. Census Bureau (2016, 2010); U.S. Department of Education (2024).
Figure 3: Demographic alignment (via share of each race) for the recommendations provided by all evaluated models across the inclusion of dialectal features [Dia.], associated entities [Ent.], and explicit indications [Exp.]. Left boxplots [ ] represents users of the same race as the figure’s calculation and right boxplots [ ] represents users of all other races. Shaded range [ ] is the range (25th - 75th percentile) of race share for baseline prompts across models.
@@ 3 RQ1: Inferred Identity Biases Recommendations
In our first experiment, we examine whether LLMs provide recommendations that are biased towards the perceived racial identity of the user, regardless of whether identity is explicitly revealed.
@@ 3.1 Demographic Alignment by Signal
A model that personalizes only when the user explicitly asks for their racial identity to factor into decisions – and otherwise avoids incorporating stereotypes – should behave differently depending on the type of identity disclosure. When identity is revealed implicitly and it is unclear whether it is intentional, we might expect the demographic alignment of recommendations to be similar to those generated when identity is not revealed. On the other hand, when identity is explicitly indicated, we expect the models to favor recommendations that have greater demographic alignment.
Figure 3 shows the degree of demographic alignment for each user identity averaged across models. We see that models respond strongly to explicit signals of user identity by providing recommendations that have comparatively high shares of the user’s race. For example, in the plot of Black users’ results for neighborhood selection, we see that when a user requests recommendations for neighborhoods and includes “keep in mind, I am Black," they receive recommended neighborhoods with around significantly more Black residents on average than a user who does not reveal their identity. This personalization based on explicit identity disclosure is seen for users of all races except White users, whose results mostly still fall squarely within the range of those given to baseline users.
Models also produce demographically aligned results when users reveal their identity implicitly, either by referencing stereotypically associated entities, or when a user writes their request in a racially-associated dialect. This occurs to a lesser extent than it does with explicitly disclosed identity, but the difference is still consistently statistically significant when the user references associated entities and occasionally significant (albeit minimal) when the user uses dialectal phrasing (evaluated though an unpaired T-Test). In the context of important life decisions, basing decisions on unintentionally disclosed racial identities can perpetuate stereotyping without the knowledge of the user. Additionally, there is the risk of error; not everyone who references an entity belongs to the stereotypically associated race (e.g., anyone can play basketball) and not everyone who speaks a specific racial dialect identifies with that race, especially with the internet offering more cross-cultural interactions Reyes (2005); Roth-Gordon et al. (2020).
@@ 3.2 Recommendation Diversity
Model Asian Black Hispanic White
GPT 4o Mini 0.51 0.48 0.47 0.51
GPT 4o 0.55 0.49 0.50 0.55
Claude 3.5 Haiku 0.50 0.52 0.48 0.53
Claude 3.5 Sonnet 0.50 0.45 0.46 0.51
Llama 3.1 70B 0.24 0.11 0.16 0.22
Llama 3.1 405B 0.42 0.30 0.35 0.46
Gemini 1.5 Pro 0.45 0.53 0.51 0.62
Reality 0.83 0.77 0.81 0.87
(a) Diversity: Normalized entropy of recommendation distributions, compared to entropy of the real-world distribution. Higher entropy indicates higher diversity
Model Asian Black Hispanic White
GPT 4o Mini 0.63 0.62 0.63 0.65
GPT 4o 0.57 0.62 0.57 0.61
Claude 3.5 Haiku 0.68 0.71 0.64 0.65
Claude 3.5 Sonnet 0.62 0.68 0.62 0.65
Llama 3.1 70B 0.83 0.88 0.83 0.83
Llama 3.1 405B 0.74 0.82 0.71 0.63
Gemini 1.5 Pro 0.59 0.51 0.53 0.54
(b) Representativenes: Normalized JSD between neighborhood recommendations and real-world distribution. Lower JSD indicates higher representativeness.
Table 1: Comparison between recommendation distributions (when race is explicitly indicated) and real-world distributions.
To understand LLMs’ ability to personalize when pragmatically indicated to do so, we also measure the diversity and representativeness of model recommendations when race is explicitly mentioned. We measure diversity as the entropy of model recommendations over the full spectrum of possible neighborhood outcomes (where complete data is more readily available), when race is explicitly indicated by the user. To avoid issues with unequal support across distributions, we use additive smoothing with $\epsilon=1\cdot 10^{-10}$. We measure representativeness as the Jensen-Shannon Divergence of the recommendation distribution and reality. We normalize both Diversity and Representativeness. Exact equations for these measurements are included in Appendix C.3
Empirically, in Table 1 we see that the models produce recommendations that are much less diverse than reality. This effect is especially pronounced for Llama 3.1 70B, where recommended neighborhoods for Black users represent just $7.8\%$ of the Black population in the cities, compared to recommendations from Gemini 1.5 Pro’s, for example, which represent over $50\%$. We also see that Llama 3.1 70B produces the recommendations that are least similar to the real-world distribution. These results indicate that when a user expresses their race in a manner that indicates personalization is desirable, the models produce less diverse and often unrepresentative results.
@@ 3.3 Stereotypical Descriptions
Beyond the numerical demographic alignment of recommended entities, we find that language models tend to bias the values they latently attribute to users of different races, represented by the variation in explanations. To examine this, we find words with the highest association with a particular demographic group. Details of this calculation along with the top twenty terms associated with each race can be found in Appendix D.
For university recommendations, we see that models include race-based terms more frequently when race is mentioned explicitly, which is pragmatically appropriate and demonstrates personalization. However, even when the only variation is dialectal (that is, all users describe the same concern around cost and similar stats), non-White students see an increased emphasis on features such as in-state resident tuition (Asian users), lenient admissions standards (Hispanic users), and opportunities for part-time work (Black users). On the other hand, when identity is explicitly disclosed, Asian users are assumed to want stem-focused programs while black users are instead told about agricultural programs.
Similarly, we see some clear patterns even among purely dialectal prompts for neighborhood recommendations. Although all users of all races express same set of transportation constraints, there is more emphasis on different modes of transportation for each racial group. Black users are told their recommendations are subway and bus-accessible, Hispanic users are tempted with bike lanes, and Asian users are promised their neighborhoods are car-accessible. Hispanic users are given neighborhoods described as chef-friendly while Asian users are given those described as foodie-friendly, reflecting stereotypes about the racial makeup of the service industry. When race is stated explicitly, the most frequently referenced stereotypes are different, but still clear; Black users are recommended more politically progressive areas than the moderate areas suggested to White users. Similarly, Hispanic and Asian users are recommended tight-knit and traditional communities, which reflect the cultural values associated with ethnic enclaves.
By generating explanations that assign different values and preferences to different racial groups, models not only produce biased recommendations, but generate rationales that reflect (and possibly reinforce) the stereotypes that already exist in American culture.
Race White Black Hispanic Asian
Acknowledgment of Race Yes No N/A Yes No N/A Yes No N/A Yes No N/A
GPT 4o Mini 78.6 78.17 29.15 11.18 11.17 32.31 11.62 10.3 7.37 4.1 3.61
GPT 4o 78.99 77.61 28.31 11.07 9.4 35.15 14.24 13.02 9.65 4.48 3.93
Claude 3.5 Haiku 83.86 79.35 78.04 24.5 7.9 7.27 30.73 17.98 14.92 8.81 4.99 4.53
Claude 3.5 Sonnet 80.58 81.11 80.54 21.74 9.17 8.49 27.43 12.82 9.7 7.05 4.29 3.92
Llama 3.1 70B 81.54 83.25 83.85 24.59 10.86 8.91 31.38 9.6 5.78 8.61 3.29 2.52
Llama 3.1 405B 82.97 83.09 37.27 9.63 9.07 44.67 10.14 5.8 15.7 6.27 2.83
Gemini 1.5 Pro 73.52 76.8 76.35 28.32 9.5 8.67 28.76 14.48 13.12 8.64 6.11 5.61
@@
Race White Black Hispanic Asian
Acknowledgment of Race Yes No N/A Yes No N/A Yes No N/A Yes No N/A
GPT 4o Mini 44.13 43.95 37.83 15.09 13.11 46.2 30.9 28.61 20.53 12.61 11.54
GPT 4o 39.3 39.22 35.62 15.89 12.46 49.03 38.28 35.23 18.45 11.94 10.6
Claude 3.5 Haiku 51.41 47.21 46.56 24.41 10.88 9.37 43.75 31.99 29.06 14.47 12.89 12.47
Claude 3.5 Sonnet 49.99 42.35 42.57 36.53 17.11 12.5 47.18 38.66 32.59 17.48 12.94 9.9
Llama 3.1 70B 7.26555Llama 3.1 70B almost always (across its very few cases) assumed the wrong race of White users, resulting in this dip. 53.13 49.83 55.24 11.68 7.18 49.3 34.7 28.03 22.88 13.54 12.1
Llama 3.1 405B 50.75 51.46 46.51 12.42 8.86 45.42 34.3 26.08 19.79 11.44 10.66
Gemini 1.5 Pro 56.37 41.06 38.96 46.79 17.82 11.28 54.74 39.6 33.52 31.89 16.43 13.93
@@
Table 2: Mean share of the user’s race across provided recommendations controlling for whether the model acknowledges the influence of race (Yes), doesn’t acknowledge the influence of race (No), and cases where no identity information was disclosed (N/A). Bolded values are significantly ($p<0.0001$) more demographically aligned than baseline recs.